# Adiabaticity Crossover: From Anderson Localization to Planckian Diffusion

### v2

@@ Abstract
We investigate electron transport in one dimension from the quantum-acoustic perspective, where the coherent-state representation of lattice vibrations results in a time-dependent deformation potential whose rate is set by the sound speed, fluctuation spectrum is set by the temperature, and overall amplitude is set by the electron-lattice coupling strength. We introduce an acceleration-based adiabatic criterion, consistent with the adiabatic theorem and Landau-Zener theory, that separates adiabatic and diabatic dynamics across the $(T,v)$ plane. The discrete classification agrees with a continuous mean-squared acceleration scale and correlates with a coherence measure given by the ratio of coherence length to the initial packet width $L_{\phi}(t)/\sigma_{0}$. We identify a broad Planckian domain in which the dimensionless diffusivity $\alpha\!=\!Dm/\hbar$ is of order unity and only weakly depends on the parameters. This domain is more prevalent in diabatic regions and in areas of reduced phase coherence, indicating a dephasing driven crossover from Anderson localization to Planckian diffusion. Using the Einstein relation together with nearly constant $\alpha$, we directly obtain a low temperature tendency $1/\tau_{\rm tr}\propto T$, offering a insight to $T$-linear resistivity in strange metals. These results provide a unified picture that links adiabaticity, dephasing, and Planckian diffusion in dynamically disordered quantum-acoustics.
@@ I Introduction
The adiabatic theorem and the Landau-Zener framework provide guidance for dynamics under slow parameter changes [Born1928, doi:10.1143/JPSJ.5.435, 10.1098/rspa.1932.0165, Landau1932_PZS_LZ, Stueckelberg1932_HPA, Majorana1932, Shevchenko2010_PhysRep_LZSReview]. Adiabatic passage across an avoided crossing follows the adiabatic eigenstate and therefore switches diabatic channels, whereas diabatic passage follows the original channel. In crystalline solids time dependence naturally arises from lattice vibrations. From the quantum-acoustic perspective, these vibrations act as a deformation potential whose rate of change in time is set by the sound speed, fluctuation spectrum is set by the temperature through thermal mode populations, and overall amplitude is set by the electron-lattice coupling strength [PhysRevB.106.054311]. Recent work has used this coherent-state picture to analyze Planckian resistivity and the role of lattice-driven dephasing in metals and model systems [doi:10.1073/pnas.2404853121, e26070552].
Standard diagnostics of adiabaticity typically evaluate instantaneous gaps and nonadiabatic couplings or invoke two-level Landau-Zener probabilities [10.1063/1.2798382]. In chemical physics, mixed quantum-classical surface hopping triggers stochastic hops between adiabatic surfaces based on coupling strength and fewest-switches logic [10.1063/1.459170]. For a multi-mode, time-dependent deformation potential, avoided crossings are dense and no single sweep rate characterizes the drive. Repeated diagonalization becomes costly and basis dependent, and classical surface assumptions obscure phase information. This motivates basis-independent observables that diagnose adiabatic or diabatic motion directly in real time.
@@
Here we develop such a framework in one dimension. We drive an electronic wave packet with a time-dependent deformation potential and introduce an acceleration-based adiabatic criterion that is consistent with the adiabatic theorem and Landau-Zener theory. We plot regime maps across temperature and sound speed, compare this discrete classification with a continuous measure from the mean-squared acceleration, and quantify residual phase coherence through the dimensionless ratio of coherence length to the initial packet width. We then connect coherence loss to transport and identify a broad Planckian domain in which the dimensionless diffusivity $\alpha\!=\!Dm/\hbar$ is of order unity [zhang2024planckiandiffusionghostanderson]. The picture that emerges is that adiabaticity tracks how much phase coherence the system retains, while diabatic dynamics correlate with dephasing that drives a crossover from Anderson localization to Planckian diffusion. We outline implications for low temperature strange metal phenomenology and motivate experiments that correlate phase-coherence measures with transport in low-dimensional materials subject to engineered dynamic disorder [doi:10.1126/science.1227612, Legros2019, Grissonnanche2021].
@@ B Acceleration-Based “Kink” Criterion of Adiabaticity
Across temperature $T$, sound speed $v$, and electron-lattice coupling strength $g$, we distinguish adiabatic and diabatic crossovers using an acceleration-based adiabatic criterion that is consistent with the adiabatic theorem and Landau-Zener theory [Born1928, doi:10.1143/JPSJ.5.435, 10.1098/rspa.1932.0165, Shevchenko2010_PhysRep_LZSReview]. Two common diagnostics in the literature are as follows. First, adiabatic-basis tests evaluate instantaneous gaps and nonadiabatic couplings and judge motion adiabatic when $\lvert\langle m|\dot{n}\rangle\rvert$ is small relative to the gap scale or when the two-level Landau-Zener probability is exponentially small [10.1063/1.2798382]. Second, mixed quantum-classical surface hopping triggers stochastic hops between adiabatic surfaces using the nonadiabatic coupling strength and fewest-switches logic [10.1063/1.459170]. In our system the deformation field is multi-mode and time fluctuating, avoided crossings are dense, there is no single sweep rate, and repeated diagonalization is costly and basis dependent, so these diagnostics become fragile. Surface hopping also presumes well separated potential energy surfaces (PESs) and classical nuclei, whereas we propagate a single quantum wave packet in a continuum-like field and must retain phase information. This motivates a basis-independent, directly observable diagnostic built from the center-of-mass acceleration.
@@ C Quantifying Phase Coherence: Coherence Length
$\mathcal{C}(\Delta x,t)=\Big\langle\big|g^{(1)}(x,x+\Delta x;t)\big|\Big\rangle_{x}$ (12)
@@
Figure 1: Acceleration-based adiabatic criterion illustrated with space-time density, center-of-mass acceleration, and equal-time snapshots. The top row shows an adiabatic case and the bottom row shows a diabatic case. Panels (c) and (f) are snapshots at $t=0.750$ ps where blue curves plot $|\psi(x,t)|^{2}$ and red curves plot the instantaneous deformation potential $V(x,t)$. Panels (a)-(c) share adiabatic parameters and panels (d)-(f) share another set of diabatic parameters. (a) Adiabatic xxtt map $|\psi(x,t)|^{2}$ with discrete kinks that mark resolved channel switches between localized pockets. Red indicates the magnitude of $|\psi|^{2}$. (b) Adiabatic $a_{\rm cm}(t)$ with many sharp accepted peaks that pass our scale-invariant peak tests. (c) Adiabatic snapshot at $t=0.750$ ps showing a compact probability profile trapped near a pocket of $V(x,t)$. (d) Diabatic xxtt map that appears smooth under our thresholds with no resolvable kinks. (e) Diabatic $a_{\rm cm}(t)$ dominated by broad undulations that fail the peak tests. (f) Diabatic snapshot at $t=0.750$ ps showing a more extended and asymmetric probability profile relative to $V(x,t)$.
@@ A Acceleration-Based Adiabatic Criterion: Adiabatic and Diabatic Dynamics
Representative space-time maps and acceleration traces in Figure 1 summarize the two dynamical sectors. In the adiabatic regime [Figure 1 (a-c)] the xxtt map shows a discrete sequence of kinks that form a clear zigzag track and the center-of-mass acceleration contains many sharp accepted peaks. The equal-time snapshot in Figure 1 (c) at $t=0.750$ ps reveals a compact probability profile aligned with a pocket of the deformation potential, consistent with channel switches between instantaneous localized eigenstates. In the diabatic regime [Figure 1 (d-f)] the xxtt map is smooth because the landscape moves too rapidly for projection onto a single instantaneous eigenstate, so no peaks pass our thresholds; the snapshot in Figure 1 (f) at $t=0.750$ ps shows a broader and more asymmetric profile. Weak incoherent fluctuations then act as an effective dephasing bath that suppresses phase-coherent backscattering and reduces the residual coherence [RevModPhys.57.287, 10.1063/1.5054017]. Empirically the diabatic sector occupies a broader portion of the $(T,v)$ map and it is not uniform. Within it lies the Planckian domain where the dimensionless diffusivity $\alpha\!=\!Dm/\hbar$ is of order unity [zhang2024planckiandiffusionghostanderson], and there are also zones toward larger sound speed $v$ that exhibit even larger diffusivity, consistent with a faster sweep of the deformation potential and stronger dephasing. All of these trajectories fail the peak tests and produce smooth acceleration traces, but they organize into bands of transport strength across the map. In both cases the long-time evolution is diffusive, which indicates that Anderson localization is disrupted to different degrees. In the example of Figure 1 the diffusivity in the diabatic case is about ten times larger than in the adiabatic case, which shows that the diabatic regime usually suppresses localization more strongly.
@@ B Phase Diagrams: Regime Map, Planckian Domain, and Quantified Coherence
Figure 2: Adiabatic-diabatic regime map in the $(T,v)$ plane at fixed $g$ from the acceleration-based adiabatic criterion. Red symbols denote adiabatic points, blue symbols denote diabatic points, and the black staircase marks the boundary.
@@
We next relate dynamical regime, phase coherence, and Planckian diffusion. Anderson localization in one dimension relies on phase-coherent backscattering [RevModPhys.57.287, PhysRevLett.55.2696]. Purely periodic motion of the landscape does not by itself erase coherence; dephasing requires fluctuations that carry effectively irreversible information [10.1063/1.5054017]. In the quantum-acoustic setting the multi-mode deformation potential has thermal phases and a colored temporal spectrum, which act as dynamic disorder and promote dephasing [PhysRevB.106.054311], related crossovers in time-fluctuating random potentials have been observed in photorefractive media [https://doi.org/10.1002/lpor.202501144]. Diabatic dynamics experience stronger exposure to these fluctuations and thus stronger dephasing, whereas adiabatic dynamics feature well-separated, resolvable channel switches with weaker dephasing. Figure 5 quantifies the remaining coherence via the dimensionless ratio $L_{\phi}/\sigma_{0}$ and shows that the contour where $L_{\phi}/\sigma_{0}\!\sim\!1$ tracks the adiabatic-diabatic divide.
@@ IV Discussion
Conceptually, Anderson localization in one dimension arises from phase-coherent backscattering [RevModPhys.57.287, PhysRevLett.55.2696]. A strictly periodic drive does not eliminate interference on its own. Dynamic disorder that carries fluctuations and randomness acts as an intrinsic dephasing environment [10.1063/1.5054017]. Our simulations realize the latter through a multi-mode acoustic field with thermal phases [PhysRevB.106.054311]. The adiabatic theorem and the Landau-Zener picture provide the operative mechanism [Born1928, 10.1098/rspa.1932.0165, Shevchenko2010_PhysRep_LZSReview]. When the sweep rate across local avoided crossings is small compared with the instantaneous gap squared, the system follows an instantaneous eigenstate and produces discrete channel switches. When the sweep is fast and carries stochastic fluctuations, the adiabatic condition is violated and the dynamics are diabatic. The state predominantly follows its diabatic branch without resolving individual avoided crossings, while random phase accumulation from the fluctuating potential dephases the interfering amplitudes and suppresses localization. The data in Figure 2 through Figure 6 show that increasing $v$ promotes nonadiabaticity by accelerating the temporal sweep of the landscape, whereas decreasing $T$ weakens fluctuations, pushes the dynamics toward the diabatic side, and reduces $L_{\phi}/\sigma_{0}$. The broad Planckian domain is then a natural consequence of dephasing that suppresses localization corrections and drives the system to a nearly quantum-limited diffusive fixed point [RevModPhys.94.041002, heller2025quantumacousticsdemystifiesstrange].

### v3

@@ Abstract
We investigate electron transport in one dimension from the quantum-acoustic perspective, where the coherent-state representation of lattice vibrations results in a time-dependent deformation potential whose rate is set by the sound speed, fluctuation spectrum is set by the temperature, and overall amplitude is set by the electron-lattice coupling strength. We introduce an acceleration-based adiabatic criterion, consistent with the adiabatic theorem and Landau–Zener theory, that separates adiabatic and diabatic dynamics across the $(T,v)$ plane. The discrete classification agrees with a continuous mean-squared acceleration scale and correlates with a coherence measure given by the ratio of coherence length to the initial packet width $L_{\phi}(t)/\sigma_{0}$. We identify a broad Planckian domain in which the dimensionless diffusivity $\alpha\!=\!Dm/\hbar$ is of order unity and only weakly depends on the parameters. This domain is more prevalent in diabatic regions and in areas of reduced phase coherence, indicating a dephasing driven crossover from Anderson localization to Planckian diffusion. Using the Einstein relation together with nearly constant $\alpha$, we directly obtain a low temperature tendency $1/\tau_{\rm tr}\propto T$, offering a insight to $T$-linear resistivity in strange metals. These results provide a unified picture that links adiabaticity, dephasing, and Planckian diffusion in dynamically disordered quantum-acoustics.
@@ I Introduction
The adiabatic theorem and the Landau–Zener framework provide guidance for dynamics under slow parameter changes [Born1928, doi:10.1143/JPSJ.5.435, 10.1098/rspa.1932.0165, Landau1932_PZS_LZ, Stueckelberg1932_HPA, Majorana1932, Shevchenko2010_PhysRep_LZSReview]. Adiabatic passage across an avoided crossing follows the adiabatic eigenstate and therefore switches diabatic channels, whereas diabatic passage follows the original channel. In crystalline solids time dependence naturally arises from lattice vibrations. From the quantum-acoustic perspective, these vibrations act as a deformation potential whose rate of change in time is set by the sound speed, fluctuation spectrum is set by the temperature through thermal mode populations, and overall amplitude is set by the electron-lattice coupling strength [PhysRevB.106.054311]. Recent work has used this coherent-state picture to analyze Planckian resistivity and the role of lattice-driven dephasing in metals and model systems [doi:10.1073/pnas.2404853121, e26070552].
Standard diagnostics of adiabaticity typically evaluate instantaneous gaps and nonadiabatic couplings or invoke two-level Landau–Zener probabilities [10.1063/1.2798382]. In chemical physics, mixed quantum-classical surface hopping triggers stochastic hops between adiabatic surfaces based on coupling strength and fewest-switches logic [10.1063/1.459170]. For a multi-mode, time-dependent deformation potential, avoided crossings are dense and no single sweep rate characterizes the drive. Repeated diagonalization becomes costly and basis dependent, and classical surface assumptions obscure phase information. This motivates basis-independent observables that diagnose adiabatic or diabatic motion directly in real time.
@@
Here we develop such a framework in one dimension. We drive an electronic wave packet with a time-dependent deformation potential and introduce an acceleration-based adiabatic criterion that is consistent with the adiabatic theorem and Landau–Zener theory. We plot regime maps across temperature and sound speed, compare this discrete classification with a continuous measure from the mean-squared acceleration, and quantify residual phase coherence through the dimensionless ratio of coherence length to the initial packet width. We then connect coherence loss to transport and identify a broad Planckian domain in which the dimensionless diffusivity $\alpha\!=\!Dm/\hbar$ is of order unity [zhang2024planckiandiffusionghostanderson]. The picture that emerges is that adiabaticity tracks how much phase coherence the system retains, while diabatic dynamics correlate with dephasing that drives a crossover from Anderson localization to Planckian diffusion. We outline implications for low temperature strange metal phenomenology and motivate experiments that correlate phase-coherence measures with transport in low-dimensional materials subject to engineered dynamic disorder [doi:10.1126/science.1227612, Legros2019, Grissonnanche2021].
@@ B Acceleration-Based “Kink” Criterion of Adiabaticity
Across temperature $T$, sound speed $v$, and electron-lattice coupling strength $g$, we distinguish adiabatic and diabatic crossovers using an acceleration-based adiabatic criterion that is consistent with the adiabatic theorem and Landau–Zener theory [Born1928, doi:10.1143/JPSJ.5.435, 10.1098/rspa.1932.0165, Shevchenko2010_PhysRep_LZSReview]. Two common diagnostics in the literature are as follows. First, adiabatic-basis tests evaluate instantaneous gaps and nonadiabatic couplings and judge motion adiabatic when $\lvert\langle m|\dot{n}\rangle\rvert$ is small relative to the gap scale or when the two-level Landau–Zener probability is exponentially small [10.1063/1.2798382]. Second, mixed quantum-classical surface hopping triggers stochastic hops between adiabatic surfaces using the nonadiabatic coupling strength and fewest-switches logic [10.1063/1.459170]. In our system the deformation field is multi-mode and time fluctuating, avoided crossings are dense, there is no single sweep rate, and repeated diagonalization is costly and basis dependent, so these diagnostics become fragile. Surface hopping also presumes well separated potential energy surfaces (PESs) and classical nuclei, whereas we propagate a single quantum wave packet in a continuum-like field and must retain phase information. This motivates a basis-independent, directly observable diagnostic built from the center-of-mass acceleration.
@@ C Quantifying Phase Coherence: Coherence Length
$\mathcal{C}(\Delta x,t)=\Big|\big\langle g^{(1)}(x,x+\Delta x;t)\big\rangle_{x}\Big|$ (12)
@@
Figure 1: Acceleration-based adiabatic criterion illustrated with space-time density, center-of-mass acceleration, and equal-time snapshots. The top row shows an adiabatic case and the bottom row shows a diabatic case. Panels (c) and (f) are snapshots at $t=0.750$ ps where blue curves plot $|\psi(x,t)|^{2}$ and red curves plot the instantaneous deformation potential $V(x,t)$. Panels (a)–(c) share adiabatic parameters and panels (d)–(f) share another set of diabatic parameters. (a) Adiabatic xxtt map $|\psi(x,t)|^{2}$ with discrete kinks that mark resolved channel switches between localized pockets. Red indicates the magnitude of $|\psi|^{2}$. (b) Adiabatic $a_{\rm cm}(t)$ with many sharp accepted peaks that pass our scale-invariant peak tests. (c) Adiabatic snapshot at $t=0.750$ ps showing a compact probability profile trapped near a pocket of $V(x,t)$. (d) Diabatic xxtt map that appears smooth under our thresholds with no resolvable kinks. (e) Diabatic $a_{\rm cm}(t)$ dominated by broad undulations that fail the peak tests. (f) Diabatic snapshot at $t=0.750$ ps showing a more extended and asymmetric probability profile relative to $V(x,t)$.
@@ A Acceleration-Based Adiabatic Criterion: Adiabatic and Diabatic Dynamics
Representative space-time maps and acceleration traces in Figure 1 summarize the two dynamical sectors. In the adiabatic regime [Figure 1 (a–c)] the xxtt map shows a discrete sequence of kinks that form a clear zigzag track and the center-of-mass acceleration contains many sharp accepted peaks. The equal-time snapshot in Figure 1 (c) at $t=0.750$ ps reveals a compact probability profile aligned with a pocket of the deformation potential, consistent with channel switches between instantaneous localized eigenstates. In the diabatic regime [Figure 1 (d–f)] the xxtt map is smooth because the landscape moves too rapidly for projection onto a single instantaneous eigenstate, so no peaks pass our thresholds; the snapshot in Figure 1 (f) at $t=0.750$ ps shows a broader and more asymmetric profile. Weak incoherent fluctuations then act as an effective dephasing bath that suppresses phase-coherent backscattering and reduces the residual coherence [RevModPhys.57.287, 10.1063/1.5054017]. Empirically the diabatic sector occupies a broader portion of the $(T,v)$ map and it is not uniform. Within it lies the Planckian domain where the dimensionless diffusivity $\alpha\!=\!Dm/\hbar$ is of order unity [zhang2024planckiandiffusionghostanderson], and there are also zones toward larger sound speed $v$ that exhibit even larger diffusivity, consistent with a faster sweep of the deformation potential and stronger dephasing. All of these trajectories fail the peak tests and produce smooth acceleration traces, but they organize into bands of transport strength across the map. In both cases the long-time evolution is diffusive, which indicates that Anderson localization is disrupted to different degrees. In the example of Figure 1 the diffusivity in the diabatic case is about ten times larger than in the adiabatic case, which shows that the diabatic regime usually suppresses localization more strongly.
@@ B Phase Diagrams: Regime Map, Planckian Domain, and Quantified Coherence
Figure 2: Adiabatic–diabatic regime map in the $(T,v)$ plane at fixed $g$ from the acceleration-based adiabatic criterion. Red symbols denote adiabatic points, blue symbols denote diabatic points, and the black staircase marks the boundary.
@@
We next relate dynamical regime, phase coherence, and Planckian diffusion. Anderson localization in one dimension relies on phase-coherent backscattering [RevModPhys.57.287, PhysRevLett.55.2696]. Purely periodic motion of the landscape does not by itself erase coherence; dephasing requires fluctuations that carry effectively irreversible information [10.1063/1.5054017]. In the quantum-acoustic setting the multi-mode deformation potential has thermal phases and a colored temporal spectrum, which act as dynamic disorder and promote dephasing [PhysRevB.106.054311], related crossovers in time-fluctuating random potentials have been observed in photorefractive media [https://doi.org/10.1002/lpor.202501144]. Diabatic dynamics experience stronger exposure to these fluctuations and thus stronger dephasing, whereas adiabatic dynamics feature well-separated, resolvable channel switches with weaker dephasing. Figure 5 quantifies the remaining coherence via the dimensionless ratio $L_{\phi}/\sigma_{0}$ and shows that the contour where $L_{\phi}/\sigma_{0}\!\sim\!1$ tracks the adiabatic–diabatic divide.
@@ IV Discussion
Conceptually, Anderson localization in one dimension arises from phase-coherent backscattering [RevModPhys.57.287, PhysRevLett.55.2696]. A strictly periodic drive does not eliminate interference on its own. Dynamic disorder that carries fluctuations and randomness acts as an intrinsic dephasing environment [10.1063/1.5054017]. Our simulations realize the latter through a multi-mode acoustic field with thermal phases [PhysRevB.106.054311]. The adiabatic theorem and the Landau–Zener picture provide the operative mechanism [Born1928, 10.1098/rspa.1932.0165, Shevchenko2010_PhysRep_LZSReview]. When the sweep rate across local avoided crossings is small compared with the instantaneous gap squared, the system follows an instantaneous eigenstate and produces discrete channel switches. When the sweep is fast and carries stochastic fluctuations, the adiabatic condition is violated and the dynamics are diabatic. The state predominantly follows its diabatic branch without resolving individual avoided crossings, while random phase accumulation from the fluctuating potential dephases the interfering amplitudes and suppresses localization. The data in Figure 2 through Figure 6 show that increasing $v$ promotes nonadiabaticity by accelerating the temporal sweep of the landscape, whereas decreasing $T$ weakens fluctuations, pushes the dynamics toward the diabatic side, and reduces $L_{\phi}/\sigma_{0}$. The broad Planckian domain is then a natural consequence of dephasing that suppresses localization corrections and drives the system to a nearly quantum-limited diffusive fixed point [RevModPhys.94.041002, heller2025quantumacousticsdemystifiesstrange].